What two numbers can be multiply to get -60 and add to -17 ?
Answer by stanbon(75887)

What two numbers can be multiply to get -60 and add to -17 ?
-------
x*y = -60
x + y = -17
-------
Substitute for "y" and solve for "x"::
x(-x-17) = -60
------
x(x+17) = 60
x^2 + 17x - 60 = 0
------
x = [-17 +- sqrt(17^2-4*-60)[]/(2)
-----
x = [-17 +- 23]/2
----
x = -20 or x = 3
------
If x = -20, y = 3
If x = 3, y = -20
----------------
Cheers,
Stan H
----------------